# PUMA: Efficient Continual Graph Learning with Graph Condensation

### v1

@@ 4.2 Pseudo Label-guided Edge-free Memory Bank
Pseudo-labelling technique is wildly used on the semi-supervised classification problem. It can enlarge the training dataset [50] and provide additional supervision signal [51]. Distribution matching involves calculating class centres of training set for the original and condensed dataset in the embedding space and aligning these centres of the condensed dataset with the counterpart of original one. The more original data points involved in the calculation, the more accurate the centre is. Therefore, unlabelled nodes with pseudo labels can be used to enlarge the training set during graph condensation.
@@
After obtaining confidence scores of pseudo labels, a threshold can be used to filter out the more certain pseudo labels for reducing the noisy labels. Distribution matching algorithm can enjoy the enlarged training set to accurately condense. The overall procedure of PUMA is shown in Algorithm 1.
@@ 8             $\tilde{E}_{c,k}=\text{MLP}_{\theta_{m}}(\boldsymbol{\tilde{X}}_{k})$
14Train a classifier $\text{GNN}_{PL}$ using $\mathcal{M}_{k-1}$ and $\mathcal{\tilde{G}}_{k}$ to generate pseudo labels $\boldsymbol{Y}_{k,PL}$ for tested/unlabelled nodes in the $\mathcal{G}_{k}$;
15 for $n\leftarrow 1$ to $N$ do
@@ 19             $\tilde{E}_{c,k}=\text{MLP}_{\theta_{n}}(\boldsymbol{\tilde{X}}_{k})$
ER-GNN only stores individual nodes, which can reduce the memory space and improve the replay efficient, but ignore the structure information. SSM keeps the structure by sampling sparsified subgraphs from the incoming graphs, which improve the model’s performance but not enjoy the aggregate-free training as ER-GNN. PUMA contains edge-free graphs which can be efficiently stored in the memory and be trained by a MLP model.
@@ 4.3 Train in Memory from Scratch
In continual learning, the vanilla replay-based CGL methods are faced with an imbalanced learning problem. When the size of the incoming graph is significantly larger than that of replayed graphs, the model is hard to balance the learning of knowledge from the historical graphs and the incoming graph. The previous attempts for balance are based on the loss scaling [8, 10], which can be represented as:
$\ell_{\text{replay}}=\alpha\mathcal{L}(\mathcal{G}_{k};\theta_{k})+\beta\mathcal{L}(\mathcal{M}_{k-1};\theta_{k}),$ (12)
where most effort is dedicated to $\alpha$ and $\beta$ according to the imbalance scale, which inevitably compromises the performance.
@@
$\begin{split}\ell_{\text{CaT}}&=\mathcal{L}(\mathcal{M}_{k};\theta_{k})\\ &=\mathcal{L}(\mathcal{\tilde{G}}_{k};\theta_{k})+\mathcal{L}(\mathcal{M}_{k-1};\theta_{k}).\end{split}$ (14)
This process is named Train in Memory (TiM) since the model only trains with replayed graphs in the memory bank.
@@
For better optimisation, the model weights of each layer are reinitialised before learning form new memories. The architecture of the CGL backbone model remains constant, such as the number of hidden layers and hidden layer dimension during the continual training process.
In summary, the proposed CaT framework uses graph condensation to generate small and effective replayed graphs and applies the TiM scheme to solve the imbalanced learning in CGL. The overall procedure of PUMA is shown in Algorithm 2.
@@ 5.1 Setup
5.1.1 Datasets
Following the previous work [11, 10, 28], four datasets for node classification tasks are used in experiments, CoraFull [52], Arxiv [53], Reddit [21] and Products [53]. CoraFull and Arxiv are both citation networks. Reddit is a post-to-post graph. The Products dataset is the co-purchase network. Table I shows the statistics of these datasets.
@@ 5.1.1 Datasets
Each dataset is split into a series of tasks focusing on the node classification problem. Each task includes nodes of two unique classes as an incoming graph. In each task, 60% nodes are chosen as training nodes, 20% nodes are for validation, and 20% are for testing. Only transductive setting is considered in this paper. Therefore, the 20% testing nodes can be observed in the graph condensation phase. Class-IL is the main focus of the experiment since it is more challenging than task-IL, although overall performance in the task-IL setting will also be reported. In the continual update phase, the model can only access the newly incoming graph and the memory bank. In the testing phase, the model is required to be evaluated with test graphs from all previous tasks. There are no inter-task edges between any two tasks. In both task-IL and class-IL setting, since the total class number is not given in advance, the output dimension is incremental as new classes appear.
@@ 5.1.2 Baselines
  • Joint is the ideal upper bound situation where the memory bank contains all historical incoming graphs.
  • EWC [22] applies quadratic penalties to the model weights that are important to the previous tasks.
  • MAS [23] utilises a regularisation term for parameters sensitive to the model performance of historical tasks.
  • GEM [24] modifies the gradients using the informative data stored in memory.
  • TWP [9] preserves the topological information for previous tasks by a regularisation term.
  • LwF[25] distils the knowledge from the old model to the new model to keep the previous knowledge.
  • HPNs [11] redesign the conventional graph embedding generation for the task-IL setting by maintaining three-level prototypes. Although HPNs are recently published, this baseline is only reported in task-IL experiments.
  • ER-GNN [8] samples the informative nodes from incoming graphs into the memory bank.
  • SSM [10] stores the sparsified incoming graph in the memory bank for future replay.
  • CaT [1] is the first and the state-of-the-art method to make use of graph condensation in continual learning, which is proposed in our previous work.
@@ 5.1.3 Evaluation Metrics
Mean of average performance (mAP) [54] denotes the average performance of model snapshots in the continual learning process:
@@ 5.1.4 Implementation
The budget ratio represents the proportion of the memory bank to the total number of nodes in the entire training set, and the budget for every task is evenly assigned. By default, the budget ratio for the Joint baseline is 1 as it stores every training data in its memory. Unless otherwise specified, for the replay-based method, the default budget ratio is 0.01, which means the size of the memory bank becomes 0.5% of the size of the entire training data. Although the budget is set to a real number instead of a ratio of the entire training set in more piratical scenarios, the budget ratio is used in the experiments for keeping fairness and comparing the efficiency of different memory banks.
In the condensation phase, a 1-layer MLP with 4096 dimensions is used to encode all datasets by default. The learning rate for updating condensed features is 0.001. In the continual training phase, a 3-layer MLP with two 512-dimensional hidden layers and a class number-dependent output layer is used as the CGL model for condensation-based methods. In contrast, other replay-based methods utilise a 3-layer GCN model, aligning with the dimensions used in the condensation-based methods. However, for methods that do not involve replay, following the experimental setup in CaT [1], only a single 256-dimensional hidden layer is used. For each incoming new task, the classifier are trained 500 epochs with 0.001 learning rate. All results are obtained by running five times and reported with the average value and the standard error, and all experiments are conducted on one NVIDIA A100 GPU (80GB).
Category Methods CoraFull Arxiv Reddit Products
AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$
Lower bound Finetuning 2.2±0.0 -96.6±0.1 5.0±0.0 -96.7±0.1 5.0±0.0 -99.6±0.0 4.3±0.0 -97.2±0.1
Regularisation EWC 2.9±0.2 -96.1±0.3 5.0±0.0 -96.8±0.1 5.3±0.6 -99.2±0.7 7.6±1.1 -91.7±1.4
MAS 2.2±0.0 -94.1±0.6 4.9±0.0 -95.0±0.7 10.7±1.4 -92.7±1.5 10.1±0.6 -89.0±0.5
GEM 2.5±0.1 -96.6±0.1 5.0±0.0 -96.8±0.1 5.3±0.5 -99.3±0.5 4.3±0.1 -96.8±0.1
TWP 21.2±3.2 -67.4±1.6 4.3±1.1 -93.0±8.3 9.5±2.0 -35.5±5.5 6.8±3.5 -64.3±12.8
Distillation LWF 2.2±0.0 -96.6±0.1 5.0±0.0 -96.8±0.1 5.0±0.0 -99.5±0.0 4.3±0.0 -96.8±0.2
Replay ER-GNN 3.1±0.2 -94.6±0.2 23.2±0.5 -77.1±0.5 20.0±3.0 -83.7±3.1 34.0±1.0 -55.7±0.8
SSM 3.5±0.5 -94.7±0.5 26.4±0.8 -73.7±0.9 41.8±3.2 -60.8±3.4 58.1±0.4 -29.3±0.5
Full dataset Joint 85.3±0.1 -2.7±0.0 63.5±0.3 -15.7±0.4 98.2±0.0 -0.5±0.0 72.2±0.4 -5.3±0.5
Condensation CaT (ours) 68.5±0.9 -5.7±1.3 64.9±0.3 -12.5±0.8 97.7±0.1 -0.4±0.1 71.1±0.3 -5.4±0.3
PUMA (ours) 77.9±0.2 -4.2±0.9 67.0±0.1 -12.2±0.3 98.0±0.1 -0.3±0.1 74.2±0.4 -4.1±0.5
TABLE III: Overall results in task-IL setting without inter-task edges. $*{}^{*}$he results of HPNs are from the original paper, and only Arxiv and Products are provided here.
Category Methods CoraFull Arxiv Reddit Products
AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$
Lower bound Finetuning 51.0±3.4 -46.2±3.5 67.1±5.2 -31.3±5.6 57.1±7.4 -44.6±7.8 56.4±3.8 -42.4±4.0
Regularisation EWC 87.4±2.2 -9.1±2.2 85.6±7.7 -11.9±8.1 85.5±3.3 -14.8±3.5 90.3±1.8 -6.8±1.9
MAS 93.0±0.3 -0.7±0.5 83.8±6.9 -12.0±7.8 99.0±0.1 0.0±0.0 95.9±0.1 0.0±0.0
GEM 94.3±0.6 -2.1±0.5 94.7±0.1 -2.3±0.2 99.3±0.1 -0.3±0.1 86.9±0.9 -10.6±0.9
TWP 87.9±1.9 -4.9±0.6 77.1±7.3 -3.5±5.4 74.1±5.5 -1.5±0.5 75.5±4.4 -4.9±6.4
Distillation LwF 64.7±1.1 -32.3±1.2 60.2±5.8 -38.6±6.2 62.4±3.5 -39.1±3.7 50.1±0.7 -49.3±0.8
Architecture $\text{HPNs}^{*}$ 85.8±0.7 0.6±0.9 - - 80.1±0.8 2.9±1.0
Replay ER-GNN 47.9±1.9 -49.1±2.0 87.3±0.8 -10.2±0.8 87.5±2.5 -12.8±2.7 80.5±1.3 -17.3±1.3
SSM 73.4±2.4 -22.8±2.5 91.8±0.4 -5.5±0.5 94.0±2.7 -5.9±2.8 92.3±0.8 -4.8±0.8
Full dataset Joint 97.2±0.0 0.2±0.1 96.7±0.0 -0.1±0.1 99.7±0.0 0.0±0.0 95.7±0.7 -0.2±0.7
Condensation CaT (ours) 93.3±0.4 -0.3±0.6 94.7±0.3 -0.8±0.3 99.3±0.0 -0.0±0.1 94.9±0.3 -0.5±0.5
PUMA (ours) 95.2±0.3 -0.7±0.2 95.3±0.1 0.1±0.1 99.4±0.0 0.0±0.0 95.4±0.3 0.1±0.5

### v2

@@ 4.2 Pseudo Label-guided Edge-free Memory Bank
Pseudo-labelling technique is widely used on the semi-supervised classification problem. It can enlarge the training dataset [11] and provide additional supervision signal [18]. Distribution matching involves calculating class centres of training set for the original and condensed dataset in the embedding space and aligning these centres of the condensed dataset with the counterpart of original one. Therefore, unlabelled nodes with pseudo labels can be used to enlarge the training set during graph condensation.
@@
After obtaining confidence scores of pseudo labels, a threshold is used to filter out the uncertain pseudo labels for reducing the noisy labels. Distribution matching algorithm can enjoy the enlarged training set to accurately condense. The overall procedure of PUMA is shown in Algorithm 1.
@@ 8             $\tilde{E}_{c,k}=\text{MLP}_{\theta_{m}}(\boldsymbol{\tilde{X}}_{k})$
14
15Train a classifier $\text{GNN}_{PL}$ using $\mathcal{M}_{k-1}$ and $\mathcal{\tilde{G}}_{k}$ to generate pseudo labels $\boldsymbol{Y}_{k,PL}$ for tested/unlabelled nodes in the $\mathcal{G}_{k}$;
16
17for $n\leftarrow 1$ to $N$ do
@@ 21             $\tilde{E}_{c,k}=\text{MLP}_{\theta_{n}}(\boldsymbol{\tilde{X}}_{k})$
ER-GNN only stores individual nodes, which can reduce the memory space and improve the replay efficiency, but ignore the structure information. SSM keeps the structure by sampling sparsified subgraphs from the incoming graphs, which improves the model’s performance but does not enjoy the aggregate-free training as ER-GNN. PUMA contains edge-free graphs which can be efficiently stored in the memory and be trained by a MLP model.
@@ 4.3 Train in Memory from Scratch
In continual learning, the vanilla replay-based CGL methods are faced with an imbalanced learning problem. When the size of the incoming graph is significantly larger than that of replayed graphs, the model is hard to balance the learning of knowledge from the historical graphs and the incoming graph. The previous attempts for balance are based on the loss scaling, $\alpha$ and $\beta$ [65, 57], which can be represented as:
$\ell_{\text{replay}}=\alpha\mathcal{L}(\mathcal{G}_{k};\theta_{k})+\beta\mathcal{L}(\mathcal{M}_{k-1};\theta_{k}).$ (12)
@@
$\begin{split}\ell_{\text{CaT}}&=\mathcal{L}(\mathcal{M}_{k};\theta_{k})\\ &=\mathcal{L}(\mathcal{\tilde{G}}_{k};\theta_{k})+\mathcal{L}(\mathcal{M}_{k-1};\theta_{k}),\end{split}$ (14)
which is called Train in Memory (TiM), since the model only trains with replayed graphs in the memory bank.
@@
For an effective optimisation, the model weights of each layer are reinitialised before learning from new memories. The architecture of the CGL backbone model remains unchanged, including the number of hidden layers and hidden layer dimensions during the continual training process.
In summary, the proposed PUMA framework uses graph condensation to generate small and effective replayed graphs, and applies the TiM scheme to solve the imbalanced learning in CGL, with details in Algorithm 2.
@@ 5.1 Setup
5.1.1 Datasets and Preprocessing
Following the previous work [57, 56], six datasets for node classification are used for evaluation, CoraFull [31], Arxiv [13], Reddit [10], Products [13], Amazon computers and Amazon photo [30]. Table I shows the dataset statistics.
@@ 5.1.1 Datasets and Preprocessing
A-Computer 13,752 491,722 767 10 5
A-Photo 7,650 238,162 745 8 4
Each dataset is split into a series of tasks. Each task includes nodes of two unique classes, which form an incoming graph. In each task, 60% nodes are chosen as training nodes, 20% nodes are for validation, and 20% are for testing. Only the transductive setting is considered in this paper. Class-IL is the main focus of the experiment, as it is more challenging than task-IL, although overall performance in both settings is reported. In the training phase, the model can only access the newly incoming graph and the memory bank. In the testing phase, the model is required to test on all existing graphs. There are no inter-task edges between any two tasks.
@@ 5.1.2 Baselines
  • Joint is the ideal upper bound situation where the memory bank contains all incoming graphs.
  • EWC (2016) [17] applies quadratic penalties to the model weights that are important to previous tasks.
  • MAS (2018) [1] utilises a regularisation for sensitive parameters to model performance on historical tasks.
  • GEM (2017) [28] modifies the gradients using the informative data stored in memory.
  • TWP (2021) [23] preserves the topological information for previous tasks by a regularisation term.
  • LwF (2018)[21] distils knowledge from the old model to the new model to keep the previous knowledge.
  • ER-GNN (2021) [65] samples the informative nodes from incoming graphs into the memory bank.
  • SSM (2022) [57] stores the sparsified incoming graph in the memory bank for future replay.
  • DSLR (2024) [4] considers both the representativeness and diversity of replayed nodes.
  • HPNs (2023) [58] redesign the embedding generation for the task-IL by maintaining three-level prototypes.
  • MetaCLGraph (2023) [39] guides model updates by memory banks and gradients from incoming graphs.
  • ER-LS-GS (2023) [12] preserves the local and the global structure of replayed graphs by regularisation.
  • CaT (2023) [26] utilises graph condensation in CGL, which is proposed in our previous work.
@@ 5.1.3 Evaluation Metrics
Mean of average performance (mAP) [64] denotes the average performance of the model in continual learning:
@@
Category Methods CoraFull Arxiv Reddit Products A-Computer A-Photo
AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$
Lower bound Finetuning 2.2±0.0 -96.6±0.1 5.0±0.0 -96.7±0.1 5.0±0.0 -99.6±0.0 4.3±0.0 -97.2±0.1 19.6±0.2 -99.2±0.2 24.2±0.1 -95.2±7.3
Regularisation EWC 2.9±0.2 -96.1±0.3 5.0±0.0 -96.8±0.1 5.3±0.6 -99.2±0.7 7.6±1.1 -91.7±1.4 19.6±0.2 -99.2±0.2 24.2±0.1 -94.5±8.1
MAS 2.2±0.0 -94.1±0.6 4.9±0.0 -95.0±0.7 10.7±1.4 -92.7±1.5 10.1±0.6 -89.0±0.5 19.6±0.1 -99.1±0.1 23.5±0.8 -98.4±0.2
GEM 2.5±0.1 -96.6±0.1 5.0±0.0 -96.8±0.1 5.3±0.5 -99.3±0.5 4.3±0.1 -96.8±0.1 19.6±0.2 -99.2±0.2 24.2±0.1 -94.5±8.1
TWP 21.2±3.2 -67.4±1.6 4.3±1.1 -93.0±8.3 9.5±2.0 -35.5±5.5 6.8±3.5 -64.3±12.8 18.9±0.9 -96.8±2.9 23.0±1.2 -98.8±0.3
Distillation LWF 2.2±0.0 -96.6±0.1 5.0±0.0 -96.8±0.1 5.0±0.0 -99.5±0.0 4.3±0.0 -96.8±0.2 19.7±0.0 -98.6±1.4 23.7±1.2 -96.9±3.7
Replay ER-GNN 3.1±0.2 -94.6±0.2 23.2±0.5 -77.1±0.5 20.0±3.0 -83.7±3.1 34.0±1.0 -55.7±0.8 19.9±0.0 -99.0±0.1 24.3±0.1 -98.7±0.1
SSM 3.5±0.5 -94.7±0.5 26.4±0.8 -73.7±0.9 41.8±3.2 -60.8±3.4 58.1±0.4 -29.3±0.5 28.6±8.6 -88.2±10.7 25.7±1.9 -96.7±2.4
DSLR 3.9±0.7 -94.2±0.4 25.2±2.5 -75.0±2.7 25.0±0.4 -51.5±0.6 37.5±1.3 -43.6±1.2 24.9±0.4 -51.7±0.7 24.8±1.0 -97.9±1.4
Ensemble MetaCLGraph 63.8±1.5 -24.7±1.6 30.9±1.1 -51.7±0.7 71.9±4.4 -28.6±4.6 19.9±0.0 -99.0±0.1 29.3±19.6 -47.1±23.5 19.1±10.5 -76.6±6.0
ER-LS-GS 3.0±0.2 -94.7±0.3 24.0±0.9 -76.3±0.9 24.9±0.2 -51.6±0.7 34.1±0.8 -55.4±0.9 19.9±0.0 -99.0±0.1 24.3±0.1 -98.6±0.1
Full dataset Joint 85.3±0.1 -2.7±0.0 63.5±0.3 -15.7±0.4 98.2±0.0 -0.5±0.0 72.2±0.4 -5.3±0.5 96.4±2.1 -1.9±2.7 97.3±0.1 -0.9±0.1
Condensation CaT (ours) 68.5±0.9 -5.7±1.3 64.9±0.3 -12.5±0.8 97.7±0.1 -0.4±0.1 71.1±0.3 -5.4±0.3 90.5±1.1 -3.1±0.7 92.0±0.6 -1.9±0.5
PUMA (ours) 77.9±0.2 -4.2±0.9 67.0±0.1 -12.2±0.3 98.0±0.1 -0.3±0.1 74.2±0.4 -4.1±0.5 91.7±0.4 -2.6±0.2 93.5±0.5 -1.8±0.5
TABLE III: Overall results in task-IL setting without inter-task edges. The results of HPNs are from the original paper.
Category Methods CoraFull Arxiv Reddit Products A-Computer A-Photo
AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$ AP (%) $\uparrow$ BWT (%) $\uparrow$
Lower bound Finetuning 51.0±3.4 -46.2±3.5 67.1±5.2 -31.3±5.6 57.1±7.4 -44.6±7.8 56.4±3.8 -42.4±4.0 48.7±12.4 -62.5±14.9 71.6±10.1 -32.0±17.2
Regularisation EWC 87.4±2.2 -9.1±2.2 85.6±7.7 -11.9±8.1 85.5±3.3 -14.8±3.5 90.3±1.8 -6.8±1.9 75.8±5.3 -29.0±6.6 74.7±7.1 -27.2±10.5
MAS 93.0±0.3 -0.7±0.5 83.8±6.9 -12.0±7.8 99.0±0.1 0.0±0.0 95.9±0.1 0.0±0.0 77.7±11.0 -26.6±13.6 82.6±11.5 -19.7±14.8
GEM 94.3±0.6 -2.1±0.5 94.7±0.1 -2.3±0.2 99.3±0.1 -0.3±0.1 86.9±0.9 -10.6±0.9 78.96±8.4 -22.59±9.6 93.05±3.5 -6.35±2.3
TWP 87.9±1.9 -4.9±0.6 77.1±7.3 -3.5±5.4 74.1±5.5 -1.5±0.5 75.5±4.4 -4.9±6.4 79.3±13.5 -23.1±16.2 86.0±7.9 -13.3±11.8
Distillation LwF 64.7±1.1 -32.3±1.2 60.2±5.8 -38.6±6.2 62.4±3.5 -39.1±3.7 50.1±0.7 -49.3±0.8 78.2±9.1 -26.2±11.3 89.1±5.0 -11.4±7.0
Architecture $\text{HPNs}^{*}$ 85.8±0.7 0.6±0.9 - - 80.1±0.8 2.9±1.0 - - - -
Replay ER-GNN 47.9±1.9 -49.1±2.0 87.3±0.8 -10.2±0.8 87.5±2.5 -12.8±2.7 80.5±1.3 -17.3±1.3 71.0±11.3 -35.2±14.1 73.6±14.0 -32.9±18.6
SSM 73.4±2.4 -22.8±2.5 91.8±0.4 -5.5±0.5 94.0±2.7 -5.9±2.8 92.3±0.8 -4.8±0.8 92.4±3.0 -8.5±3.7 95.2±1.3 -4.1±1.6
DSLR 71.7±3.3 -24.5±3.2 91.8±0.4 -5.6±0.4 49.8±1.5 -1.9±1.9 48.3±0.8 -4.9±1.3 79.4±11.5 -24.7±14.4 92.3±5.0 -8.0±6.4
Ensemble MetaCLGraph 92.5±0.9 -10.0±0.9 84.8±1.1 -5.5±1.1 91.3±1.7 -3.2±1.1 92.5±2.3 -4.1±0.9 80.0±16.3 -11.9±12.1 69.1±16.7 -18.2±22.6
ER-LS-GS 47.4±2.3 -49.5±2.4 87.7±0.8 -9.9±0.8 52.2±1.3 -3.0±3.3 80.3±1.5 -17.4±1.4 68.4±9.9 -38.4±12.4 62.4±7.0 -47.9±9.3
Full dataset Joint 97.2±0.0 0.2±0.1 96.7±0.0 -0.1±0.1 99.7±0.0 0.0±0.0 95.7±0.7 -0.2±0.7 98.9±0.1 -0.0±0.1 98.3±0.0 0.1±0.1
Condensation CaT (ours) 93.3±0.4 -0.3±0.6 94.7±0.3 -0.8±0.3 99.3±0.0 -0.0±0.1 94.9±0.3 -0.5±0.5 97.7±0.2 -0.2±0.2 96.5±1.5 0.4±0.4
PUMA (ours) 95.2±0.3 -0.7±0.2 95.3±0.1 0.1±0.1 99.4±0.0 0.0±0.0 95.4±0.3 0.1±0.5 97.8±0.1 -0.1±0.1 97.3±0.2 -0.2±0.2
@@ 5.1.4 Implementation
The budget ratio represents the proportion of the memory bank to the total number of nodes in the entire training set, and the budget for every task is evenly assigned. By default, the budget ratio for the Joint baseline is 1 as it stores every training data in its memory. Unless otherwise specified, for the replay-based method, the default budget ratio is 0.005, which means the size of the memory bank becomes 0.5% of the size of the entire training data. Although the budget is set to a real number instead of a ratio of the entire training set in more piratical scenarios, the budget ratio is used in the experiments for keeping fairness and comparing the efficiency of different memory banks.
In the condensation phase, a 1-layer MLP with 4096 dimensions is used to encode all datasets by default. The learning rate for updating condensed features is 0.001. In the continual training phase, a 3-layer MLP with two 512-dimensional hidden layers and a class number-dependent output layer is used as the CGL model for condensation-based methods. In contrast, other replay-based methods utilise a 3-layer GCN with 512 hidden dimensions. All results are obtained by running five times and reported with the average value and the standard error, and all experiments are conducted on one NVIDIA A100 GPU (80GB).